OpenStudy (anonymous):

if sinx+sin^2x=1 then what is the value of cos^12x+3cos^10x+3cos^8x+cos^6x

OpenStudy (anonymous):

\[(a+b)^{3}=a^{3}+3a^{2}b+3ab^{2}+b^{3}\] and \[\cos^{2}(x)+\sin^{2}(x)=1\] May be useful

OpenStudy (anonymous):

sin(x)+sin^2(x)=1 so sin(x)=cos^2(x) (by cos^2(x)+sin^2(x)=1) then cos^12(x)+3cos^10(x)+3cos^8(x)+cos^6(x) can be written as sin^6(x)+3sin^5(x)+3sin^4(x)+sin^3(x) =(sin^2(x)+sin(x))^3 =1^3 =1